OpenStudy (anonymous):

@kelliegirl33 I do not want an answer! Lucy rode on a train that traveled at a constant rate. The graph compares the number of hours she traveled to the number of miles she traveled. What was the rate of speed of the train? A. 50 miles per hour B. 75 miles per hour <--- C. 100 miles per hour D. 125 miles per hour

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (acxbox22):

you are correct because every hour the train travels an additional 75 miles in the graph 75 is the slope

OpenStudy (candyme):

What you need to do is set up and equation in y=mx+b form. Your b is the starting speed, m is the rate that you are going to find, and y and x are points on your line. Plug your b in and then plug other points to find the m

OpenStudy (anonymous):

speed = distance/time speed = 150/2 = 75 or speed = 300/4 = 75 no matter what point you pick, the average speed wil be 75 mph

OpenStudy (anonymous):

I didn't give you the amswer last time...I just told you that I agreed with your answer and showed you how I did it